OpenStudy (anonymous):

Andrea Gilford's savings account has a balance of $169. After 2 years, the amount of interest at 12% compounded quarterly will be $ . C. The stories told by los viejitos inspired Anaya's love of words and books.

OpenStudy (tkhunny):

Have you considered calculating it? Annual Interest is 12% Quarterly Interest is 12% / 4 = 3% Go!

OpenStudy (wolf1728):

12% Quarterly interest = [(1 + r/n)^1/n]-1 Where 'n' is number of compounding periods. (in this case it's 4) [(1 + .12/4) ^ 4]-1 =1.12550881 -1 =12.550881 annual interest

OpenStudy (tkhunny):

Why in the name of reason do we want the annual interest rate? Let's just solve the problem directly. \(169 - 169(1.03)^{8} =\;??\)

OpenStudy (tkhunny):

Whoops! Other way around.

OpenStudy (wolf1728):

total amount = principal * (1 +r)^n total amount = 169 * (1.12550881)^2 total amount = 169*1.2667700814 total amount = 214.08 calculator to check your work: http://www.1728.org/compint.htm

OpenStudy (wolf1728):

We want the annual rate because that is what the quarterly compounding will yield.

OpenStudy (wolf1728):

If we use the formula total = 169 * (1+r/n)^n2 where n is the number of compounding periods per year and n2 is years * n, we get the same answer $214.08 However we have to remember with this formula, we need to use 2 values of 'n'. total = 169 * (1+12/4)^8